# Emoji Reactions on Telegram Often Reflect Social Approval Over Emotional Resonance

### v1

@@ Abstract
Emoji reactions are a frequently used feature of messaging platforms. Prior work mainly interpreted emojis as indicators of emotional resonance or user sentiment. However, emoji reactions may instead reflect broader social dynamics. Here, we investigate the communicative function of emoji reactions on Telegram by analyzing the relationship between the emotional and rhetorical content of messages and the emoji reactions they receive. We collect and analyze over 650k Telegram messages that received at least one emoji reaction. We annotate each message with sentiment, emotion, persuasion strategy, and speech act labels, and infer the sentiment and emotion of emoji reactions using both lexicons and large languages. We find a systematic mismatch between message sentiment and reaction sentiment, with positive reactions dominating even when the message is neutral or negative. We show that this pattern remains consistent across rhetorical strategies and emotional tones, suggesting that emoji reactions may signal a degree of social approval rather than reflecting emotional resonance. Finally, we shed light on the communicative strategies that predict greater emoji engagement. These findings have methodological implications for sentiment analysis, as interpreting emoji reactions as direct proxies for emotional response may be misleading.
Emoji Reactions on Telegram
Often Reflect Social Approval Over Emotional Resonance
Serena Tardelli1, Lorenzo Alvisi1,2, Lorenzo Cima1,3, Stefano Cresci1, Maurizio Tesconi1 1Institute of Informatics and Telematics, National Research Council, Italy 2IMT School for Advanced Studies Lucca, Italy 3University of Pisa, Italy {firstname.lastname}@iit.cnr.it
@@ 1 Introduction
Emoji reactions are a central feature of digital platforms, enabling users to respond quickly and visually to content. While much research has focused on in-text emojis (i.e., linguistic elements embedded within authored messages), emoji reactions (i.e., reactions appended by readers as lightweight feedback mechanisms), remain comparatively underexplored. Yet they are equally revealing, as they are instantaneous, costless, and a quick way to interact with content without composing a reply. Moreover, while prior work has examined in-text emojis, showing they amplify sentiment Shiha and Ayvaz (2017), improve sentiment classification Liu et al. (2021), and provide emotional nuance Lou et al. (2024); Khan et al. (2025), this understanding has often been extended implicitly to emoji reactions Pool and Nissim (2016), treating them as affective signals. However, this assumption has rarely been tested empirically. Recent work suggests that emoji reactions may instead express pragmatic or social meanings, such as to approve or normalize toxic discourse Morales et al. (2025).
Here we investigate the communicative function of emoji reactions on Telegram, a messaging platform widely used for news, politics, and activism. On Telegram, emoji reactions are publicly visible and frequently used, making them a rich signal of collective response. In detail, we analyze over 650k Telegram messages and their associated emoji reactions. We combine large-scale emotion classification with rhetorical analysis to examine the alignment between message tone and reaction type. In particular, we address three research questions:
@@
  • RQ2: Do emoji reactions reflect emotional resonance?
  • RQ3: What message features predict emoji reactions?
@@
  • We provide the first large-scale analysis of emoji reactions alignment on Telegram to identify which reactions distort or enhance emotional alignment.
  • We show a consistent misalignment between sentiment of the message and emoji reactions.
  • We show that emoji reactions do not always reflect emotional resonance with the message, but also act as signals of social support.
  • We identify specific message features, such as emotional framing or persuasive tactics, that predict user engagement.
Together, our findings challenge the assumption that emoji reactions straightforwardly reflect emotional response, particularly in socially dynamic platforms like Telegram. Instead, they reveal a richer set of communicative functions, where emoji reactions often act as signals of social approval, encouragement, or group affiliation.
@@ 2 Related work
On a similar line, recent work Morales et al. (2025) explored how emoji reactions are used in toxic versus non-toxic contexts on Telegram, showing that even toxic comments often receive positive emoji reactions, signaling approval or normalization of harmful content. Our findings complement and generalize this pattern: while Morales et al. (2025) focused specifically on toxicity, we observe that positive reaction bias holds across a broader range of communicative intentions, including persuasive tactics, emotional tone, and speech acts. Moreover, while our results do not exclude the possibility of emotional mirroring, they suggest that social signaling frequently overrides it in practice.
@@
By focusing on Telegram, a relatively understudied platform despite its growing relevance in political discourse and information diffusion, we contribute new evidence to this emerging body of work. While recent studies examined phenomena such as toxicity and misinformation on Telegram Urman and Katz (2022); Alvisi et al. (2024, 2025), much less is known about the communicative function of emoji reactions in these environments.
Our analysis suggests that emoji reactions are not reliable proxies for emotional content, and that their interpretation requires careful attention to platform norms, community practices, and the communicative function of approval. This has important implications for sentiment modeling, misinformation research, and theories of social signaling in digital environments.
@@ 3 Dataset
We collect our dataset by retrieving Telegram channels from TGStat [^1] https://tgstat.com/, a third-party analytics platform that categorizes public channels by topic, language, and popularity. We select only English-language channels with at least 5,000 subscribers, spanning a wide range of thematic categories. This initial sampling yields 20,695 channels. For each channel, we collect all public messages posted between January and December 2024, and focus on those that received at least one emoji reaction. The resulting dataset comprises 647,879 messages from 993 chats, all related to the topic of cryptocurrencies, as emerged from the data. Message distribution follows a long-tail pattern consistent with a power law, with few channels accounting for most messages. While Telegram supports both default and custom emojis for reactions, our analysis includes only standard emojis.
@@ 4 Methods
We analyze Telegram messages using a multi-stage pipeline that combines NLP-based linguistic annotation with behavioral cues derived from user reaction data (i.e., emoji-based reactions). First, we annotate each message with linguistic features, including sentiment, persuasion strategies, emotion, and speech act categories. Then, we annotate the emoji reactions to infer both their sentiment polarity and emotional content. This allows capturing both linguistic content and social responses.
@@ 5.2 RQ2 - Emoji reactions and emotional resonance
While RQ1 revealed a tendency for messages to receive positive emoji reactions regardless of their sentiment, the analysis was based on a coarse-grained polarity framework (positive, neutral, negative). To better understand which emotions drive this misalignment, RQ2 shifts focus from sentiment to discrete emotion categories. We ask whether certain emoji reactions systematically distort alignment with the emotional tone of the message, and whether removing them improves interpretability. Specifically, we hypothesize that emoji reactions are not always used to mirror the emotional tone of a message, but instead often serve as signals of social approval or engagement. To evaluate this, we measure how often the emotion of the message appears among the top-ranked emoji reactions. In other words, we compute the percentage of messages where the top-ranked emoji reaction shares the same emotion category as the message itself. To assess the impact of individual emoji reactions, we perform an ablation analysis. We remove one emotion at a time from the reaction distribution and we re-calculate the top-ranked emoji reaction. If removing a specific emotion significantly increases the number of matches (i.e., if it makes the dominant emoji reaction more likely to align with the emotion of the message), we interpret that category as disruptive to emotional alignment (i.e., the removed reaction was often selected even when it did not match the emotional tone of the message). Figure 3 presents the relative change in alignment match after removing each emoji reaction category, grouped by the emotion label of the message. Each cell shows the percentage change in the proportion of messages where the emotion of the top-ranked emoji matches the emotion of the message. Statistically significant changes, defined as those with a standardized effect size (Cohen’s h) $\geq 0.2$, are marked with an asterisk. Cohen’s h quantifies the magnitude of difference between proportions before and after ablation, providing an interpretable measure of effect strength beyond percentages.
@@
The consistent improvement in alignment when joy reactions are removed suggests that these emojis are often used even when they do not match the emotional tone of the message. Unlike other categories, joy appears to systematically distort emotional alignment, pointing to a distinct functional role. Rather than reflecting shared affect, joy reactions may serve pragmatic or social purposes, such as signaling encouragement, approval, or support, regardless of the emotional content of the message. The disruptive effect of joy, revealed through targeted ablation, highlights its unique communicative function in digital contexts, one that is orthogonal to emotional resonance.
@@ 6 Discussion and Conclusions
Our analysis reveals that emoji reactions on Telegram often reflect social signaling rather than emotional resonance. While most prior work has focused on emojis embedded in text, showing that they amplify sentiment Shiha and Ayvaz (2017), enhance sentiment classification Liu et al. (2021), and provide valuable emotional context Lou et al. (2024); Khan et al. (2025), this view has also been implicitly extended to emoji reactions, treating them as proxies for sentiment Pool and Nissim (2016). Our findings challenge this assumption. Through ablation analysis, we show that joy is the most disruptive emotion in terms of alignment. Unlike other categories, joy reactions frequently appear in response to emotionally negative or neutral messages. This suggests that joy serves a pragmatic role, signaling approval, support, or social bonding. These patterns indicate that emoji reactions on Telegram may operate more as social cues than as emotional reflections. This urges caution against using emoji reactions as ground truth for sentiment analysis, particularly in environments like Telegram where social signaling may override emotional congruence. It also raises concerns for AI models that might estimate public opinion based on emoji reactions Our results may also help explain why negative or controversial content online often receives higher engagement, as reactions may reflect support for the person posting. These findings also complement work by Morales et al. (2025), where authors highlighted positive emoji reactions in the context of toxic discourse as approval signal. Similarly, we show that this bias toward positive reactions holds across a wider range of communicative intents and emotional tones. This is also consistent with the nature of Telegram groups and channels, where users voluntarily join communities that may align with their views. Finally, we show that stronger negative sentiment and muted positivity are linked to higher emoji reaction rates, extending the negativity bias observed in online platforms Schöne et al. (2023); Watson et al. (2024) to emoji-based interactions.
In conclusion, we provided insights into the social meaning of emoji reactions on Telegram, suggesting that they serve a distinct communicative purpose compared to in-text emojis. Recognizing this distinction is especially important for research on online engagement, misinformation, political discourse, and human–computer interaction, where understanding the why behind reactions is as important as measuring their presence. We invite future work to account for the different roles these emojis play, to more effectively capture the pragmatic and social dimensions of emoji use on messaging platforms, informing approaches to communication strategy, content moderation, and platform design.
@@ 7 Limitations
This study has several limitations. First, our findings rely on automated classifiers, such as sentiment and emotion models based on BERT or gpt-4o, which have not been validated on our dataset. Second, our dataset consists mostly of public Telegram channels related to cryptocurrency, which may influence the communicative dynamics we observe, as these might reflect social norms specific to that domain. In addition, the distribution of activity is uneven, as some large channels contribute higher number of messages, potentially skewing the results toward dominant community patterns. Finally, the Telegram reaction system itself introduces some structural constraints. The set of standard emoji reactions is finite, and our study does not account for custom emoji sets. Moreover, not all channels allow reactions at all, and when they do, administrators can restrict the available emojis to a small, predominantly positive set. This can introduce selection bias and encourage approval-oriented responses and positive signaling by design.

### v2

# Emoji Reactions on Telegram: Unreliable Indicators of Emotional Resonance

@@ Abstract
Emoji reactions are a frequently used feature of messaging platforms, yet their communicative role remains understudied. Prior work on emojis has focused predominantly on in-text usage, showing that emojis embedded in messages tend to amplify and mirror the author’s affective tone. This evidence has often been extended to emoji reactions, treating them as indicators of emotional resonance or user sentiment. However, they may reflect broader social dynamics. Here, we investigate the communicative function of emoji reactions on Telegram. We analyze over 650k crypto-related messages that received at least one reaction, annotating each with sentiment, emotion, persuasion strategy, and speech act labels, and inferring the sentiment and emotion of emoji reactions using both lexicons and LLMs. We uncover a systematic mismatch between message and reaction sentiment, with positive reactions dominating even for neutral or negative content. This pattern persists across rhetorical strategies and emotional tones, indicating that emojis used as reactions do not reliably function as indicators of emotional mirroring or resonance of the content, in contrast to findings reported for in-text emojis. Finally, we identify the features that most predict emoji engagement. Overall, our findings caution against treating emoji reactions as sentiment labels, highlighting the need for more nuanced approaches in sentiment and engagement analysis.
Emoji Reactions on Telegram:
Unreliable Indicators of Emotional Resonance
Serena Tardelli1, Lorenzo Alvisi1,2, Lorenzo Cima1,3, Stefano Cresci1, Maurizio Tesconi1 1Institute of Informatics and Telematics, National Research Council, Italy 2IMT School for Advanced Studies Lucca, Italy 3Department of Information Engineering, University of Pisa, Italy {firstname.lastname}@iit.cnr.it
@@ 1 Introduction
Emoji reactions are a central feature of digital platforms, enabling users to respond quickly and visually to content. While much research has focused on in-text emojis (i.e., linguistic elements embedded within authored messages), emoji reactions (i.e., reactions appended by readers as lightweight feedback mechanisms), remain comparatively underexplored. Yet they are equally revealing, as they are instantaneous, costless, and a quick way to interact with content without composing a reply. Moreover, while prior work has examined in-text emojis, showing they amplify sentiment Shiha and Ayvaz (2017), improve sentiment classification Liu et al. (2021), and provide emotional nuance Lou et al. (2024); Khan et al. (2025), this understanding has often been extended implicitly to emoji reactions Pool and Nissim (2016), treating them as affective signals. However, recent work suggests that emoji reactions may instead express pragmatic or social meanings, such as to approve or normalize toxic discourse Morales et al. (2025).
Building on this idea, we provide the first large-scale study examining whether emoji reactions actually reflect emotional resonance, an assumption often implicitly adopted when reactions are used as proxy labels in sentiment analysis. To do so, we investigate the communicative function of emoji reactions on Telegram, a messaging platform widely used for news, politics, and activism. On Telegram, emoji reactions are publicly visible and frequently used, making them a rich signal of collective response. In detail, we analyze over 650k crypto-related Telegram messages and their associated emoji reactions. We combine large-scale emotion classification with rhetorical analysis to examine the alignment between message tone and reaction type. In particular, we address three research questions:
@@
  • RQ2: Do emoji reactions reflect the emotional tone of the original message, creating emotional resonance?
  • RQ3: Which message features predict emoji reactions?
@@
  • We provide the first large-scale analysis of emoji reactions on Telegram to determine whether they distort or amplify emotional alignment.
  • We conduct this analysis on a large dataset of 650k crypto-related Telegram channels, a domain known for strong community identity and high engagement, where reactions are especially frequent and publicly visible.
  • We show a consistent misalignment between message sentiment and emoji reactions, indicating that, unlike in-text emojis, reactions do not reliably act as emotional mirroring.
  • We identify specific message features, such as emotional framing or persuasive tactics, that predict user engagement in terms of emoji reactions.
Our findings suggest that emoji reactions are not reliable proxies for sentiment content, as they weakly mirror or resonate with the emotion evoked by the message. Together, these results challenge the assumption that emoji reactions reliably reflect emotional response, as commonly observed for in-text emojis. These patterns should also be interpreted in light of Telegram’s specific communicative context, where community norms may shape reaction behavior. More broadly, these findings have important implications, particularly for misinformation and political communication research, where reaction-based signals are often interpreted as indicators of emotional resonance or public sentiment.
@@ 2 Related work
On a similar line, recent work Morales et al. (2025) explored how emoji reactions are used in toxic versus non-toxic contexts on Telegram, showing that even toxic comments often receive positive emoji reactions, signaling approval or normalization of harmful content Cima et al. (2025). Our findings complement and generalize this pattern: while Morales et al. (2025) focused specifically on toxicity, we observe that positive reaction bias holds across a broader range of communicative intentions, including persuasive tactics, emotional tone, and speech acts. Moreover, while our results do not exclude the possibility of emotional mirroring, they suggest that social signaling frequently overrides it in practice.
@@
By focusing on Telegram, which remains comparatively understudied despite its growing role in political discourse and information diffusion, we contribute new evidence on communication dynamics within this platform. While recent studies have examined toxicity, misinformation, and content moderation on Telegram Urman and Katz (2022); Alvisi et al. (2024, 2025), the communicative function of emoji reactions remains largely unexplored, a gap that we address in the present study.
@@ 3 Dataset
We collect our dataset by retrieving Telegram channels from TGStat [^1] https://tgstat.com/, a third-party analytics platform that categorizes public channels by topic, language, and popularity. We select only English-language channels with at least 5,000 subscribers. This initial sampling yields 20,695 channels. For each channel, we collect all public messages posted between January and December 2024, and focus on those that received at least one emoji reaction. The resulting dataset comprises 647,879 messages from 993 chats related to the topic of cryptocurrencies. Message distribution follows a long-tail pattern consistent with a power law, with few channels accounting for most messages. While Telegram supports both default and custom emojis for reactions, our analysis includes only standard emojis.
@@ 4 Methods
We analyze Telegram messages using a multi-stage pipeline that combines NLP-based linguistic annotation with behavioral cues derived from user reaction data (i.e., emoji-based reactions). First, we automatically annotate each message with linguistic features, including sentiment, persuasion strategies, emotion, and speech act categories. Then, we annotate the emoji reactions to infer both their sentiment polarity and emotional content. This allows capturing both linguistic content and social responses.
@@ 5.2 RQ2 - Emoji reactions and emotional resonance
While RQ1 revealed a tendency for messages to receive positive emoji reactions regardless of their sentiment, the analysis was based on a coarse-grained polarity framework (positive, neutral, negative). To better understand which emotions drive this misalignment, RQ2 shifts focus from sentiment to discrete emotion categories. We ask whether certain emoji reactions systematically distort alignment with the emotional tone of the message, and whether removing them improves interpretability. Specifically, we hypothesize that emoji reactions are not always used to mirror the emotional tone of a message, but instead often serve as signals of social approval, engagement, or banter. To evaluate this, we measure how often the emotion of the message appears among the top-ranked emoji reactions. In other words, we compute the percentage of messages where the top-ranked emoji reaction shares the same emotion category as the message itself. To assess the impact of individual emoji reactions, we perform an ablation analysis. We remove one emotion at a time from the reaction distribution and we re-calculate the top-ranked emoji reaction. If removing a specific emotion significantly increases the number of matches (i.e., if it makes the dominant emoji reaction more likely to align with the emotion of the message), we interpret that category as disruptive to emotional alignment (i.e., the removed reaction was often selected even when it did not match the emotional tone of the message). Figure 3 presents the relative change in alignment match after removing each emoji reaction category, grouped by the emotion label of the message. Each cell shows the percentage change in the proportion of messages where the emotion of the top-ranked emoji matches the emotion of the message. Statistically significant changes, defined as those with a standardized effect size (Cohen’s h) $\geq 0.2$, are marked with an asterisk. Cohen’s h quantifies the magnitude of difference between proportions before and after ablation, providing an interpretable measure of effect strength beyond percentages.
@@
The consistent improvement in alignment when joy reactions are removed suggests that these emojis are often used even when they do not match the emotional tone of the message. Unlike other categories, joy appears to systematically distort emotional alignment, pointing to a distinct functional role. Rather than reflecting shared affect, joy reactions may serve pragmatic or social purposes, such as signaling encouragement, approval, banter, or support, regardless of the emotional content of the message. The disruptive effect of joy, revealed through targeted ablation, highlights its unique communicative function in digital contexts, one that is orthogonal to emotional resonance.
@@ 6 Discussion and Conclusions
Our analysis reveals that emoji reactions on Telegram do not reliably act as emotional mirroring and may instead signal social support. While most prior work has focused on emojis embedded in text, showing that they amplify sentiment Shiha and Ayvaz (2017), enhance sentiment classification Liu et al. (2021), and provide valuable emotional context Lou et al. (2024); Khan et al. (2025), this view has also been implicitly extended to emoji reactions, treating them as proxies for sentiment Pool and Nissim (2016). Our findings challenge this assumption. Through ablation analysis, we show that joy is the most disruptive emotion in terms of alignment. Unlike other categories, joy reactions frequently appear in response to emotionally negative or neutral messages. This suggests that joy serves a pragmatic role, signaling approval, support, or social bonding. These patterns indicate that emoji reactions on Telegram may operate more as social cues than as emotional reflections. This urges caution against using emoji reactions as ground truth for sentiment analysis. This also raises concerns for AI models that might estimate public opinion based on emoji reactions. Our results also help explain why negative or controversial content online often receives higher engagement, as reactions may reflect support for the person posting, complementing work by Morales et al. (2025), where authors highlighted positive emoji reactions in the context of toxic discourse as approval signal. Moreover, we show that this bias toward positive reactions holds across a wider range of communicative intents and emotional tones. Finally, we show that stronger negative sentiment and muted positivity are linked to higher emoji reaction rates, extending the negativity bias observed in online platforms Schöne et al. (2023); Watson et al. (2024) to emoji-based interactions. As our analysis focuses on crypto-related messages, a domain where emoji reactions are dense, public, and central to community interaction, the findings describe how emoji reactions function in this specific communicative context, while also pointing to the need to examine whether the same patterns appear in other Telegram communities or online platforms. In conclusion, we provide insights into the social meaning of emoji reactions on Telegram, suggesting that they serve a distinct communicative purpose compared to in-text emojis. Recognizing this distinction is especially important for research on online engagement, misinformation, political discourse, where emoji reactions can be used as signals or indicators of emotional resonance or public sentiment.
@@ 7 Limitations
First, our findings rely on automated classifiers, including sentiment and emotion models based on BERT or gpt-4o, which have not been specifically validated on our dataset and may therefore introduce the usual degree of uncertainty associated with automated annotations. Second, as our dataset consists mostly of public cryptocurrency-related Telegram channels, the communicative dynamics we observe may reflect domain-specific norms, which limits the generalizability of our findings and highlights the need to examine other contexts. Doing so would also clarify our interpretation of the positivity bias: although we view it as consistent with social-approval, engagement, or banter signaling, other explanations remain possible, including community norms specific to this domain that may encourage more positive reactions. Nonetheless, evidence that emoji reactions do not reliably reflect emotional resonance, at least in this setting, remains informative. At the same time, our analytical approach is general and can be applied to other domains and platforms. Beyond these interpretative factors, the Telegram reaction system itself introduces structural constraints. The set of standard emoji reactions is finite, and our study does not consider custom emoji sets. In addition, not all channels enable reactions, and when they do, administrators may restrict the available options to a small and often predominantly positive set. These features can introduce selection bias and promote approval-oriented responses by design. Future work could extend these insights by examining how reaction emojis operate within a broader range of digital communication practices.